OpenStudy (anonymous):

6. What is the gravitational potential energy of the tennis ball at the top of each hill? Hill A: 80 cm Hill B: 60 CM Hill C: 40 cm Convert to meters please

OpenStudy (stormfire1):

Conversions: Hill A = 0.80 m Hill B = 0.60 m Hill C = 0.40 m Now you need the mass of the tennis ball since gravitational potential energy (near the Earth's surface is:\[U=mgh\]Since you didn't post the mass to use let's go with 57g or 0.057 kg. Now just plug in and solve each for U: \[Ua = 0.057 kg * 9.8 m/s^2 * 0.80m = 0.44688 => ~0.45 J\]\[Ub = 0.057 kg * 9.8 m/s^2 * 0.60m = 0.33516=> ~0.34 J\]\[Uc = 0.057 kg * 9.8 m/s^2 * 0.40m = 0.22344=> ~0.22 J\]